# Weak-Jamming Detection in IEEE 802.11 Networks: Techniques, Scenarios and Mobility

### v1

@@ 3 Related Work
Table 1: Qualitative comparison of relevant literature on jamming detection. Duplicated rows with the same reference indicate multiple proposals by the same paper. For evaluation setup, ● indicates Indoor and Outdoor evaluation, ◐ indicates Indoor or Outdoor, and ○ indicates simulation or analytical evaluation. The symbol - indicates that the paper provides no information about a specific feature.
@@
[24] Packet Drop Probability SVM IEEE 802.11 OFDM AWGN
@@
In this paper, we focus on the detection of weak jamming, i.e., detecting jamming in a situation where the level of the received jammer power at a target is not enough to shut down the communication link entirely but only to affect it slightly. In our earlier contributions, we denoted such a scenario as a Low-BER Regime since it translates into a low (non-zero) BER at the target. We notice that weak-jamming detection is relevant for static and mobile networks. Mobile devices, such as with Unmanned Aerial Vehicles and autonomous vehicles, could detect jamming while moving toward the jammer’s location. In these contexts, weak-jamming detection enables the pilot and the Ground Control Station (GCS) to keep control of the vehicle rather than losing control and relying on a pre-programmed action to navigate safely. Although some works in the literature consider tests for jamming under mobility, e.g., [12] and [18], to the best of our knowledge, only a few papers delved into the detection of weak and low-power jamming and proposed solutions to this problem. The pioneering contribution by Garnaev et al. in [23] only provided numerical analysis, while the contributions by Boganoski et al. [22] and Venkata et al. [24] used simulations to prove the feasibility of detecting low-power jamming. More recently, Villain et al. [25] first used real-world data collected in an office environment to show that, using Principal Component Analysis (PCA), it is possible to detect low-power jamming. However, they do not consider the impact of the distance between transmitter, receiver and jammer, and not even mobile deployments. They also consider an external device specifically meant to jamming detection. A step further in this direction is offered by our recent contributions in [3] and [4], which cover many different properties of low-power-jamming detection, including the impact of the distance between the jammer and the communication link. They are also the only contributions releasing the data and the code used for their experiments, encouraging reproducibility. Although using real-world data, those contributions consider simple communication links, not aligning with real-world operational environments, and test their solutions indoors. In this paper, we advance further the state of the art on weak-jamming detection by (i) extending the methodology defined in [3] and [4] for more complex modulation schemes; (ii) using an actual communication technology used by many wireless deployments, including drones and connected vehicles, i.e., WiFi, (iii) evaluating weak-jamming detection with various modulation schemes used in the real-world WiFi, i.e., OFDM with BPSK, QPSK, 16-QAM and 64-QAM, and (iv) performing extensive tests outdoors, considering various signal propagation conditions (indoors, outdoor garden and outdoor field). Finally, we highlight that the image generation technique used in our paper shares similarities with the one recently used for Radio Frequency Fingerprinting (RFF), e.g., in [30][31], and [32]. The proposal in [30] is meant to detect spoofing attacks to satellite constellation through the analysis of the physical-layer information available at the receiver. To this aim, the authors generated images from IQ samples and applied classification to detect spoofing attacks. The authors in [31] showed that the power-cycles of the radios affect the performance of RFF solutions, and that techniques based on images can improve performance compared to analying raw physical-layer data. The authors in [32] show that Adversarial Machine Learning techniques can be used to improve the robustness of image-based RFF techniques, so avoiding synthetic attacks via imitation of the features of the physical-layer signals at a receiver. However, using such image generation technique for the detection of weak jamming entails different choices and analysis methods, not affecting the novelty and aim of our contribution.
@@ 5.1 Image Generation
We collect IQ samples using a device capable of acquiring PHY-layer wireless channel information, e.g., an SDR. Next, we divide these samples into subsets of $n$ samples, denoted as the number of samples per image (see Sec. 7 for details on the impact of this parameter on jamming detection). Then, we represent each subset of $n$ IQ samples through the IQ constellation diagram, with the In-phase component on the x-axis and the Quadrature component on the y-axis. We highlight that our image processing technique differs from the previously proposed approach, being more suitable for higher-order modulation schemes. Previous solutions [3] and [4] considered only one of the two clouds in the image. In contrast, we consider the IQ constellation with x-axis and y-axis limits of $[-2,2]$ and divide it into a $P\times Q$ grid. Note that, at the receiver, we normalize the value of the IQ samples, thus ensuring that their values do not fall outside of the considered range. We use such a grid to generate a bi-variate histogram by counting how many IQ samples fall into a specific tile of the grid. The bins in the histogram are considered to be the value of the image pixels. Lastly, the output of the histogram is saved as a $P\times Q$ grayscale image, where $P$ and $Q$ are selected based on the input requirements of the classifier used later in the jamming detection phase. In Sec. 7, we experimentally compare the proposed image generation method with the technique used in [3] and [4] and evaluate the pros and cons.
@@ 5.2 Jamming Detection
We consider a training process characterized by 30 epochs and a mini-batch size of 35 elements. We use images generated from IQ samples captured with no jamming affecting the communication link (NO JAM images) and an equal number of images generated from IQ samples captured with jamming (referred to as JAM images) for the training process. The final output of the algorithm is the decision between jamming or no jamming.
@@
Finally, note that we select the same values for the hyperparameters of the AE as the approach presented in [4], i.e., a hidden size value of $16$, a sparsity regularization term of $0.5$, an L2-regularization term of $0.01$, the Logistic Sigmoid (logsig) function as the encoder transfer function and $250$ epochs for the training process.
@@ 6.2 Measurement Settings
Modulation Type Jam BPSK, QPSK, 16-QAM, 64-QAM 0 79,79 0,$0$
@@ 7.1 Performance Metrics
We use accuracy as the main performance metric, computed as $\frac{TP+TN}{TP+FP+TN+FN}$, being $TP$ the number of true positives (JAM images correctly classified), $TN$ the true negatives (NO JAM images correctly classified), $FP$ the number of false positives (NO JAM images classified as JAM) and $FN$ the number of false negatives (JAM images classified as NO JAM). Although such data are available, we do not report precision and recall results since we are interested in comparing the overall performance of the investigated approaches rather than the contribution of False Positive and False Negative. To obtain our results, we use K-fold cross-validation with $K=5$ and, for each value, we report the mean and $95$% confidence intervals, computed using Matlab’s $tinv$ function, i.e., the inverse cumulative distribution function of the Student’s distribution. The confidence intervals are represented as red error bars in the bar plots, with a different color for each result in the line plots. To ensure fairness of comparison between CNNs and AEs, we used the same amount of images for training both classifiers, i.e. $150$ NO JAM images. Then, for each experiment, the testing set is constituted by an equal number of NO JAM and JAM images, depending the amount of jamming data available for the particular experiment. This number can be obtained from Tab. 3, dividing the number of IQ samples for each environment and testing conditions by the number of samples per image used. The remaining part of the NO JAM dataset has been used for validation.
For some plots, we introduce a custom metric, i.e., the Signal to Noise Ratio (SNR) Degradation Ratio, typically presented as the second x-axis. The SNR Degradation Ratio models the impact of jamming on the communication channel, and it is computed as the ratio between the average SNR of the JAM images and the average SNR of the NO JAM images, i.e., $SNR_{DR}=\frac{SNR_{JAM}}{SNR_{NOJAM}}$. Note that, due to different factors affecting SNR, the SNR Degradation ratio is not comparable across different experiments. Finally, whenever not specified, we use only measurements characterized by a BER $<0.01$, so to enforce a low-BER scenario following the definition in Sec. 2.
@@ 7.2 Image Generation Parameters
Exp. 3: Image Augmentation. We compare the accuracy of our proposed solutions with the literature, considering the adoption of various image augmentation techniques. For this test, we consider the data collected in the environment E2, as it provides a more challenging classification task, allowing us to test our solution in the worst-case scenario. We empirically select 150 images from the no-jamming dataset and 75 from the jamming dataset and investigate the effect of image augmentation against the benchmark proposals in [3] and [4] with $5\cdot 10^{4}$ samples per image. We consider the following augmentation strategies, empirically chosen from those made available by Matlab [42]: (i) 180 degrees rotation, (ii) left-right mirror, (iii) up-down mirror, (iv) contrast increase and (v) brightness increase. We report all our results in Fig. 13. Note that we report with the legend No Augmentation the performances obtained when only real-world data are used in the training set, providing a benchmark for the performance obtained using the augmentation strategies.
@@ 7.4 Communication Parameters
Exp. 9: Modulation Scheme. We consider different modulation schemes for the legitimate communication link among the supported options in IEEE 802.11g, i.e., BPSK, QPSK, 16-QAM and 64-QAM. We consider the setup shown earlier in Fig. 10, with the TX, RX, and Jammer gain values set to $66$, $50$, and $16$, respectively. Note that we chose the setup and gain values such that the BER for the modulation 64-QAM—the most complex modulation scheme with the least margin for error—is as low as possible. To adapt the image generation method proposed in the literature to higher-order modulation schemes, we change the x-axis and y-axis limits to center the created clouds while keeping the same rationale. For BPSK, we use $[0,2]$ and $[-1,1]$; for QPSK, we use $[-0.293,1.707]$ and $[-1,1]$; for 16-QAM, we use $[-0.867,2.133]$ and $[-1.5,1.5]$, and finally, for 64-QAM, we use $[-0.883,2.117]$ and $[-1.5,1.5]$. Figure 19 shows the results of our experiments, considering both jamming detection approaches and the reference solutions in [3] and [4], in the environment E2.
@@ 8 Discussion
Jamming Detection Approaches. Our results demonstrate that both AEs and the chosen CNN (Resnet-18) are reliable for weak-jamming detection. The two approaches exhibit similar performance when considering most jamming and communication parameters. However, the AE-based technique only requires not-jammed samples for the training process, i.e., the samples acquired during regular communications, which are easier to collect in practice. The CNN-based approach exhibits better performance when considering some image generation parameters (samples per image and training set size) and some boundary conditions, e.g., when the RSS of the legitimate communication link becomes comparable to the one of the jammer at the receiver side. However, being a binary classification approach, it requires training on samples from both classes (jammed and not-jammed IQ samples), which could be challenging to collect in many real-world scenarios.
@@
Limited Acquisition Bandwidth. For our experiments, the use of the LimeSDR USB forced us to use an acquisition bandwidth of $5$ MHz instead of $20$ MHz, as required by IEEE 802.11g. We emphasize that a reduced sample rate is a worst-case scenario for jamming detection at the PHY layer since less information is available to the receiver for signal classification. Thus, we believe using a higher acquisition bandwidth does not affect our findings.
Limited Experiments Area. We acknowledge that the maximum distance considered for our experiments, i.e., 23 meters, may not be sufficiently representative of extensive jamming in large outdoor scenarios. This limitation of our study is mainly due to two factors. First, the outdoor areas available for our real-world experiments is limited, and we cannot legally perform extensive jamming outdoor without affecting the quality of WiFi services available to other users. Second, the equipment used for the experiments includes SDRs, which are not professional jammers capable of disrupting communications for hundreds of meters. Such devices have a limited transmission power and, in turn, they can jam nearby communications in a limited space.
Nonetheless, we first highlight that the maximum distance considered in our experiments (23 meters) is higher than the only contribution in the current literature investigating the impact of the jammer distance on the detection accuracy, i.e., the contribution in [4], which considered a maximum distance of $21$ meters, only in an indoor environment.
Moreover, although the limitations mentioned above, we remark that such limitations do not affect the validity of our findings and analysis. In fact, the aim of our analysis is to investigate the effectiveness of detecting weak jamming using various techniques proposed in the literature, and to find the boundary conditions where such techniques could detect weak jamming. As shown in Fig. 17 and Fig. 18, when the jammer is located 23 meters away from the receiver, all the investigated jamming detection approaches are characterized by sub-optimal performances, indicating that they cannot detect weak jamming reliably. As highlighted through the upper x-axis in those figures, these performances occur with a value of the SNR degradation ratio of 1.0 for the environment E1 (Fig. 17) and 0.984 for the environment E2 (Fig. 18). Similar considerations apply for Fig. 22. Therefore, the usage of more powerful devices over larger distances does not provide additional insight to our analysis and does not affect the validity of our conclusions.
@@
Known Distance TX-RX. Finally, we note that all the investigated approaches, including the ones we compared, could require knowledge of the distance between the transmitter and receiver. Concerning the mobile scenario introduced in Sec. 4, mobile vehicles usually know such a distance from the communication link with the remote transmitter, e.g., through the adopted telemetry protocols. We recall that our solutions detect jamming early, i.e., when the jamming on the receiver is weak and the communication link is still active. Thus, information availability is never a problem, contrary to ex-post approaches in the literature.

### v2

# arXiv:2505.19633v2[cs.CR] 27 Feb 2026

@@ 3 Related Work
Table 1: Qualitative comparison of relevant literature on jamming detection. Duplicated rows with the same reference indicate multiple proposals by the same paper. For evaluation setup, ● indicates Indoor and Outdoor evaluation, ◐ indicates Indoor or Outdoor, and ○ indicates simulation or analytical evaluation. The symbol - indicates that the paper provides no information about a specific feature. The works [28][2], and [29] are not reported since, although applying image generation from RF signals, they do not deal with jamming detection.
@@
[40] Packet Drop Probability SVM, Random Forests IEEE 802.11 OFDM AWGN
@@
In this paper, we focus on the detection of weak jamming, i.e., detecting jamming in a situation where the level of the received jammer power at a target is not enough to shut down the communication link entirely but only to affect it slightly. In our earlier contributions, we denoted such a scenario as a Low-BER Regime since it translates into a low (non-zero) BER at the target. We notice that weak-jamming detection is relevant for static and mobile networks. Mobile devices, such as with Unmanned Aerial Vehicles and autonomous vehicles, could detect jamming while moving toward the jammer’s location. In these contexts, weak-jamming detection enables the pilot and the Ground Control Station (GCS) to keep control of the vehicle rather than losing control and relying on a pre-programmed action to navigate safely. Although some works in the literature consider tests for jamming under mobility, e.g., [31] and [37], to the best of our knowledge, only a few papers delved into the detection of weak and low-power jamming and proposed solutions to this problem. The pioneering contribution by Garnaev et al. in [12] only provided numerical analysis, while the contributions by Boganoski et al. [6] and Venkata et al. [40] used simulations to prove the feasibility of detecting low-power jamming. The authors in [40] used SVM and Random Forest classifiers to detect jamming, identifying early the potential of ML to address this problem. However, they considered the scenario of vehicular networks and their proposed methods took significant time (in the order of seconds), not being suitable to highly dynamic conditions. More recently, Villain et al. [41] were the first to use real-world data collected in an office environment to show that, using Principal Component Analysis (PCA), it is possible to detect low-power jamming. Such works mostly focus on spectral differences, while providing limited quantitative results across various jamming detection parameters. Indeed, they do not consider the impact of the distance between transmitter, receiver and jammer, and not even mobile deployments. They also consider an external device specifically meant for jamming detection. A step further in this direction is offered by our recent contributions in [1] and [35], which cover many different properties of low-power-jamming detection, including the impact of the distance between the jammer and the communication link. They are also the only contributions releasing the data and the code used for their experiments, encouraging reproducibility. Although using real-world data, those contributions consider simple communication links, not aligning with real-world operational environments, and test their solutions indoors. In this paper, we advance further the state of the art on weak-jamming detection by (i) extending the methodology defined in [1] and [35] for more complex modulation schemes; (ii) using an actual communication technology used by many wireless deployments, including drones and connected vehicles, i.e., WiFi, (iii) evaluating weak-jamming detection with various modulation schemes used in the real-world WiFi, i.e., OFDM with BPSK, QPSK, 16-QAM and 64-QAM, and (iv) performing extensive tests outdoors, considering various signal propagation conditions (indoors, outdoor garden and outdoor field). Moreover, compared to early works on low-power jamming detection, this paper makes significant advancements: (i) extensive real-world data collection, (ii) extensive experiments and results considering various jamming detection parameters, (iii) use of state-of-the-art DL-based solutions, and (iv) very limited collection and inference time, applicable to real-time scenarios. Finally, we highlight that the image generation technique used in our paper shares similarities with the one recently used for Radio Frequency Fingerprinting (RFF), e.g., in [28][2], and [29]. The proposal in [28] is meant to detect spoofing attacks to satellite constellation through the analysis of the physical-layer information available at the receiver. To this aim, the authors generated images from IQ samples and applied classification to detect spoofing attacks. The authors in [2] showed that the power-cycles of the radios affect the performance of RFF solutions, and that techniques based on images can improve performance compared to analying raw physical-layer data. The authors in [29] show that Adversarial Machine Learning techniques can be used to improve the robustness of image-based RFF techniques, so avoiding synthetic attacks via imitation of the features of the physical-layer signals at a receiver. However, using such image generation technique for the detection of weak jamming entails different choices and analysis methods, not affecting the novelty and aim of our contribution.
@@ 5.1 Image Generation
We collect IQ samples using a device capable of acquiring PHY-layer wireless channel information, e.g., an SDR. Next, we divide these samples into subsets of $n$ samples, denoted as the number of samples per image (see Sec. 7 for details on the impact of this parameter on jamming detection). Then, we represent each subset of $n$ IQ samples through the IQ constellation diagram, with the In-phase component on the x-axis and the Quadrature component on the y-axis. We highlight that our image processing technique differs from the previously proposed approach, being more suitable for higher-order modulation schemes. Previous solutions [1] and [35] considered only one of the two clouds in the image. In contrast, we consider the IQ constellation with x-axis and y-axis limits of [2,2][-2,2] and divide it into a $P\times Q$ grid. Note that, at the receiver, we normalize the value of the IQ samples, thus ensuring that their values do not fall outside of the considered range. We apply dataset-level normalization, i.e., we compute the maximum values of raw I and Q ($I_{MAX}$, $Q_{MAX}$), and we normalize the samples as $I=\frac{I}{I_{MAX}}$ and $Q=\frac{Q}{Q_{MAX}}$. We use the mentioned grid to generate a bi-variate histogram by counting how many IQ samples fall into a specific tile of the grid. The bins in the histogram are considered to be the value of the image pixels. Lastly, the output of the histogram is saved as a $P\times Q$ grayscale image, where $P$ and $Q$ are selected based on the input requirements of the classifier used later in the jamming detection phase. In Sec. 7, we experimentally compare the proposed image generation method with the technique used in [1] and [35] and evaluate the pros and cons.
@@ 5.2 Jamming Detection
We consider a training process characterized by 30 epochs and a mini-batch size of 35 elements. We use images generated from IQ samples captured with no jamming affecting the communication link (NO JAM images) and an equal number of images generated from IQ samples captured with jamming (referred to as JAM images) for the training process. The final output of the algorithm is the decision between jamming or no jamming. For hyper-parameters tuning, we referred to the same configuration used in [35].
@@
Finally, note that we select the same values for the hyperparameters of the AE and CNNs as the approach presented in [35], i.e., a hidden size value of $16$, a sparsity regularization term of $0.5$, an L2-regularization term of $0.01$, the Logistic Sigmoid (logsig) function as the encoder transfer function and $250$ epochs for the training process.
@@ 6.2 Measurement Settings
Modulation Type Jam BPSK, QPSK, 16-QAM, 64-QAM 0 79,79 0,$00$
@@ 7.1 Performance Metrics
We use accuracy as the main performance metric, computed as $\frac{TP+TN}{TP+FP+TN+FN}$, being $TP$ the number of true positives (JAM images correctly classified), $TN$ the true negatives (NO JAM images correctly classified), $FP$ the number of false positives (NO JAM images classified as JAM) and $FN$ the number of false negatives (JAM images classified as NO JAM). Although such data are available, we do not report precision and recall results since we are interested in comparing the overall performance of the investigated approaches rather than the contribution of False Positive and False Negative. To obtain our results, we use K-fold cross-validation with $K=5$ and, for each value, we report the mean and 9595% confidence intervals, computed using Matlab’s tinvtinv function, i.e., the inverse cumulative distribution function of the Student’s distribution. The confidence intervals are represented as red error bars in the bar plots, with a different color for each result in the line plots. To ensure fairness of comparison between CNNs and AEs, we used the same amount of images for training both classifiers, i.e. $150$ NO JAM images. Then, for each experiment, the testing set is constituted by an equal number of NO JAM and JAM images, depending the amount of jamming data available for the particular experiment. This number can be obtained from Tab. 3, dividing the number of IQ samples for each environment and testing conditions by the number of samples per image used. The remaining part of the NO JAM dataset has been used for validation.
For some plots, we introduce a custom metric, i.e., the Signal to Noise Ratio (SNR) Degradation Ratio, typically presented as the second x-axis. The SNR Degradation Ratio models the impact of jamming on the communication channel, and it is computed as the ratio between the average SNR of the JAM images and the average SNR of the NO JAM images, i.e., $SNR_{DR}=\frac{SNR_{JAM}}{SNR_{NOJAM}}$. Note that, due to different factors affecting SNR, the SNR Degradation ratio is not comparable across different experiments. Finally, whenever not specified, we use only measurements characterized by a BER <0.01<0.01, so to enforce a low-BER scenario following the definition in Sec. 2.
@@ 7.2 Image Generation Parameters
Exp. 3: Image Augmentation. We compare the accuracy of our proposed solutions with the literature, considering the adoption of various image augmentation techniques. For this test, we consider the data collected in the environment E2, as it provides a more challenging classification task, allowing us to test our solution in the worst-case scenario. We empirically select 150 images from the no-jamming dataset and 75 from the jamming dataset and investigate the effect of image augmentation against the benchmark proposals in [1] and [35] with $5\cdot 10^{4}$ samples per image. We consider the following augmentation strategies, empirically chosen from those made available by Matlab [14]: (i) 180 degrees rotation, (ii) left-right mirror, (iii) up-down mirror, (iv) contrast increase and (v) brightness increase. We report all our results in Fig. 13. We report with the legend No Augmentation the performances obtained when only real-world data are used in the training set, providing a reference benchmark.
@@ 7.4 Communication Parameters
Exp. 9: Modulation Scheme. We consider different modulation schemes for the legitimate communication link among the supported options in IEEE 802.11g, i.e., BPSK, QPSK, 16-QAM and 64-QAM. We consider the setup shown earlier in Fig. 10, with the TX, RX, and Jammer gain values set to $66$, $50$, and $16$, respectively. Note that we chose the setup and gain values such that the BER for the modulation 64-QAM—the most complex modulation scheme with the least margin for error—is as low as possible. To adapt the image generation method proposed in the literature to higher-order modulation schemes, we change the x-axis and y-axis limits to center the created clouds while keeping the same rationale. For BPSK, we use $[0,2]$ and [1,1][-1,1]; for QPSK, we use [0.293,1.707][-0.293,1.707] and [1,1][-1,1]; for 16-QAM, we use [0.867,2.133][-0.867,2.133] and [1.5,1.5][-1.5,1.5], and finally, for 64-QAM, we use [0.883,2.117][-0.883,2.117] and [1.5,1.5][-1.5,1.5]. Figure 19 shows the results of our experiments, considering both jamming detection approaches and the reference solutions in [1] and [35], in the environment E2.
@@ 8 Discussion
Jamming Detection Approaches. Our results demonstrate that both AEs and the chosen CNN (Resnet-18) are reliable for weak-jamming detection. The two approaches exhibit similar performance when considering most jamming and communication parameters. However, the AE-based technique only requires not-jammed samples for the training process, i.e., the samples acquired during regular communications, which are easier to collect in practice. The CNN-based approach exhibits better performance when considering some image generation parameters (samples per image and training set size) and some boundary conditions, e.g., when the RSS of the legitimate communication link becomes comparable to the one of the jammer at the receiver side. However, being a binary classification approach, it requires training on samples from both classes (jammed and not-jammed IQ samples), which could be challenging to collect in many real-world scenarios. Our experiments considered several measurement parameters potentially impacting jamming detection. Many of such parameters are independent, but some of them are correlated: the jammer distance, transmitter gain, and receiver distance are all correlated through the SNR at the receiver. By comparing the results for these experiments among them, it is clear that the SNR is the (direct or indirect) variable mostly affecting the capability of detecting weak-jamming: when the SNR degrades significantly or is too low, all investigated jamming detection approaches struggle to detect jamming, while performances improve with lower SNR degradation ratios or higher absolute SNR values.
Comparison across Measurement Environments. The primary aim of this research is to compare various jamming detection methods in specific environments. In general, each environment is characterized by different propagation conditions, which require different fine-tuning of the transmission power and jammer power necessary to achieve weak jamming. In turn, such different configurations make systematic comparison among environments hard to achieve in a fair way across all experiments. However, as a by-product, our extensive experimental analysis also allows limited comparison of particular methods across scenarios. For example, the results for Experiment 4 (Jammer Oversampling Ratio), Experiment 6 (Jammer Signal Type) and Experiment 7 (Jammer Location) demonstrate that jamming detection across various Jamming Oversampling Ratio (JOR) values, jammer hardware and type is always possible independently of environment type. We also notice that multipath makes jamming detection harder, especially in low-SNR conditions. This is evident by comparing Fig. 17 with Fig. 18 and Fig. 20 (a) with Fig. 20 (b): for both experiments, the presence of multipath leads to more unstable performances, indicating that multipath plays a role in making jamming detection more challenging. Future work could take these results as a starting point to execute a more systematic comparison of performances across measurement environments.
@@
Limited Acquisition Bandwidth. For our experiments, the use of the LimeSDR USB forced us to use an acquisition bandwidth of $5$ MHz instead of $20$ MHz, as required by IEEE 802.11g. This reduced sample rate is a worst-case scenario for jamming detection at the PHY layer, since less information is available to the receiver for signal classification. Thus, using larger acquisition bandwidth does not affect our findings.
Limited Experiments Area. We acknowledge that the maximum distance considered for our experiments, i.e., 23 meters, may not be sufficiently representative of extensive jamming in large outdoor scenarios. This limitation of our study is mainly due to two factors. First, the outdoor areas available for our real-world experiments is limited, and we cannot legally perform extensive jamming outdoor without affecting the quality of WiFi services available to other users. Second, the equipment used for the experiments includes SDRs, which are not professional jammers capable of disrupting communications for hundreds of meters. Such devices have a limited transmission power and, in turn, they can jam nearby communications in a limited space. On the one hand, such a limitation motivates the need for additional work examining larger jamming distances, which could enhance further the practical relevance and completeness of our study. On the other hand, we highlight that the maximum distance considered in our experiments (23 meters) is higher than the only contribution in the current literature investigating the impact of the jammer distance on the detection accuracy, i.e., the contribution in [35], which considered a maximum distance of $21$ meters, only in an indoor environment.
Moreover, although the existence of the limitations mentioned above, we remark that such limitations do not affect the validity of our findings and analysis. In fact, the aim of our analysis is to investigate the effectiveness of detecting weak jamming using various techniques proposed in the literature, and to find the boundary conditions where such techniques could detect weak jamming. As shown in Fig. 17 and Fig. 18, when the jammer is located 23 meters away from the receiver, all the investigated jamming detection approaches are characterized by sub-optimal performances, indicating that they cannot detect weak jamming reliably. As highlighted through the upper x-axis in those figures, these performances occur with a value of the SNR degradation ratio of 1.0 for the environment E1 (Fig. 17) and 0.984 for the environment E2 (Fig. 18). Similar considerations apply for Fig. 22. Therefore, the usage of more powerful devices over larger distances does not provide additional insight to our analysis and does not affect the validity of our conclusions.
Limited Training Data. For our analysis, we use up to $150$ images of the NO JAM dataset for training and an equal number of JAM images, depending on the amount of jamming data available for the particular experiment. On the one hand, the usage of such a reduced number of images is an added value of our results. On the other hand, the number of images used for training is relatively limited. Thus, practitioners aiming to deploy our solution(s) in the wild should extend such a training dataset considering domain-specific datasets, collected directly in the environment where the solution is intended to be applied.
@@
Known Distance TX-RX. All the investigated approaches could require knowledge of the distance between the transmitter and receiver. Concerning the mobile scenario introduced in Sec. 4, mobile vehicles usually know such a distance from the communication link with the remote transmitter, e.g., through the adopted telemetry protocols. We recall that our solutions detect jamming early, i.e., when the jamming on the receiver is weak and the communication link is still active. Thus, information availability is never a problem, contrary to ex-post approaches in the literature.
Scaling to Multi-channel and Multi-hop 802.11 Deployments. When considering extensive 802.11 deployments encompassing multiple channels and several hops separating IEEE 802.11 nodes, additional challenges apply. To name a few, neighboring 802.11 links could share partially overlapping channels, requiring extensive training to distinguish jamming from unintentional interference, even more with modulation schemes that admit interference by design (e.g., Non-Orthogonal Multiple Access — NOMA). Moreover, such complex networks require coordination mechanisms to detect and communicate jamming effects at scale. Thus, future research could investigate how to scale up this research to such scenarios.